OpenStudy (anonymous):

Curves in Spaaaace! Find the unit tangent vector nd the unit normal vector for the parameterized curve R(t) =

OpenStudy (anonymous):

\[T = v(t)/mag(v(t))\] \[T = <2t, 1>/\sqrt(4t^2 +1)\] \[N =(T)\prime\]

OpenStudy (anonymous):

@